OpenStudy (anonymous):

(t+9)^2/t+8 * (t+8)^2/t+9 mulitply and simplify

OpenStudy (whpalmer4):

\[\frac{(t+9)^2}{(t+8)}*\frac{(t+8)^2}{(t+9)}\] See anything you can cancel to simplify before you multiply everything out?

OpenStudy (anonymous):

not really im so lost

OpenStudy (whpalmer4):

What if I rewrite it like this: \[\frac{(t+9)^2}{(t+9)}*\frac{(t+8)^2}{(t+8)}\]

OpenStudy (anonymous):

would those be 1?

OpenStudy (whpalmer4):

Here, let me rewrite it yet again :-) \[\frac{(t+9)(t+9)}{(t+9)}*\frac{(t+8)(t+8)}{(t+8)}\]

OpenStudy (anonymous):

idk that just made me a little more lost lol sorry

OpenStudy (anonymous):

I haven't done math since high school

OpenStudy (whpalmer4):

Okay, do you agree that \[(t+9)^2 = (t+9)(t+9)\]?

OpenStudy (anonymous):

yes i get that one

OpenStudy (whpalmer4):

Good. So the last thing I did was I just replaced \( (t+9)^2\) with \((t+9)(t+9)\) and \( (t+8)^2\) with \((t+8)(t+8)\) — are we cool with that?

OpenStudy (anonymous):

that wont change the answer??

OpenStudy (whpalmer4):

no, they are equal! (something)^2 is just (something)(something)

OpenStudy (anonymous):

ok whats next

OpenStudy (whpalmer4):

And in the first step, all I did was switch the denominators around — that's okay because they are getting multiplied together, and a*b = b*a, right?

OpenStudy (anonymous):

ok

OpenStudy (whpalmer4):

so, no reason to be confused :-) Now, what can we do with a fraction if the same thing appears in the numerator and the denominator?

OpenStudy (whpalmer4):

Take for example \[\frac{2*3}{3*3}\]

OpenStudy (whpalmer4):

Do you remember canceling matching things in fractions?

OpenStudy (whpalmer4):

\[\frac{3}{3} = 1\] right?

OpenStudy (anonymous):

not at all =(

OpenStudy (anonymous):

i remember that

OpenStudy (whpalmer4):

Okay, so our fraction could be written as\[\frac{2*3}{3*3} = \frac{2}{3}*\frac{3}{3} = \frac{2}{3}*1 = \frac{2}{3}\]Okay with that?

OpenStudy (anonymous):

ok

OpenStudy (whpalmer4):

Great. That's all we need to make this problem much easier :-) \[\frac{(t+9)(t+9)}{(t+9)} * \frac{(t+8)(t+8)}{(t+8)} = \frac{(t+9)}{(t+9)}*(t+9)*\frac{(t+8)}{(t+8)}*(t+8)\]

OpenStudy (whpalmer4):

What does \[\frac{(t+9)}{(t+9)}\] equal?

OpenStudy (anonymous):

1

OpenStudy (whpalmer4):

Excellent! And the same goes for (t+8)/(t+8). What do we have left after we turn those two fractions in 1s?

OpenStudy (whpalmer4):

into 1s, that is

OpenStudy (whpalmer4):

Doesn't it become \[1*(t+9)*1*(t+8) = (t+9)(t+8)\]?

OpenStudy (whpalmer4):

Do you remember how to expand that product?

OpenStudy (anonymous):

not at all ugh

OpenStudy (whpalmer4):

Okay, it's straightforward. If you've got (a+b)(c+d), that is equal by the distributive property to: \[(a+b)(c+d) = a(c+d) + b(c+d) = ac + ad + bc + bd\] Basically, each term of the first thing gets multiplied by each term of the second thing Your problem is\[(t+9)(t+8) = \] I'll do a similar one as an example, in case the symbols don't make sense to you: \[(x+3)(x+4) = x(x+4) + 3(x+4) = x*x + 4*x + 3*x + 3*4 = \]\[x^2 + 4x + 3x + 12 = x^2 + 7x + 12\] Can you do the same procedure with yours? I'll check your answer...

OpenStudy (anonymous):

t(t+8)+9(t+8)

OpenStudy (whpalmer4):

Okay, that's the first step. What do you get after you expand that?

OpenStudy (anonymous):

i get losttttt

OpenStudy (whpalmer4):

\[t(t+8)+9(t+8) = t*t + t*8 + 9(t+8)\]Can you do the other half?

OpenStudy (anonymous):

t^2+9t+8t...

OpenStudy (whpalmer4):

+ 9*8 \[t(t+8) + 9(t+8) = t*t + t*8 + 9*t + 9*8 = t^2 + 17t + 72\] That's your final answer. Here's a nice video on multiplying binomials: https://www.khanacademy.org/math/algebra/polynomials/multiplying_polynomials/v/multiplying-binomials There are many good videos on that site, and it would probably be worth your time to watch some and brush up on your skills. I have to leave now, but I'll leave you with a different way of thinking of this multiplication problem: Another way to think of it is as if you are doing multiplication of multiple-digit numbers: t + 9 X t + 8 ------- 8t +72 <- this is 8 * (t+9) t^2 + 9t <- this is t * (t+9) ------------- t^2+17t+72

OpenStudy (anonymous):

is that in lowest terms??

OpenStudy (whpalmer4):

By the way, aren't you glad we canceled things out first? Otherwise, we would have had \[\frac{t^4+34t^3+433t^2+2448t+5184}{t^2+17t+72}\] :-) Yes, \[t^2+17t+72\] is in lowest terms, and cannot be further simplified. We don't have any fractions, we don't have any parentheses, and we only have one term with each power of t.

OpenStudy (anonymous):

its telling me that is wrong

OpenStudy (whpalmer4):

Well, maybe you copied the problem wrong...

OpenStudy (anonymous):

... no i did it exactly

OpenStudy (anonymous):

i appreciate you sticking with me though

OpenStudy (anonymous):

is that factored form??

OpenStudy (whpalmer4):

See the first choice under alternate forms.

OpenStudy (whpalmer4):

Anyhow, I have to go now. Have a look at those videos, the lecturer is usually pretty clear. Feel free to shoot me a message if you get stuck on something and I'll try to help you out the next time I'm online.

OpenStudy (anonymous):

ok thanks